Consider the following information:

Portfolio Expected Return Standard
Deviation
Risk-free 6 % 0 %
Market 10.8 24
A 8.8 13


a. Calculate the Sharpe ratios for the market portfolio and portfolio A. (Round your answers to 2 decimal places.)




b. If the simple CAPM is valid, is the above situation possible?

  • Yes

  • No
Answer #1

a. sharpe ratio of market portfolio = (expected return of portfolio - risk free return on portfolio)/ standard deviation

= (10.8 - 6)/24 = 0.2

  

sharpe ratio of market portfolio = (expected return of portfolio - risk free return on portfolio)/ standard deviation

= (8.8 - 6)/13 = 0.2154

b. yes.